OpenStudy (anonymous):

SOLVE:-3a-5b=-17 3a+8b=5

OpenStudy (ryan123345):

elimination or substitution?

OpenStudy (anonymous):

Are you solving for A or B? Or does A or B already have a value? You are not being very clear about the problem.

OpenStudy (anonymous):

rearranging equation 2 gives 3a=5-8b substituting in eqyation 1 gives -(5-8b)-5b=-17 simplifying to -5+8b-5b=-17 add 5 to both sides and combine b's to get 3b=-12 divide both sides by 3 to get b=-4 substitute -4 in either equation for b and solve to get a=37/3